OpenStudy (anonymous):

use the distributive property to write -3(7-11) as an equivalent expression

OpenStudy (anonymous):

First find out what the number is inside of the parenthesis. 7 - 11 = 4 now you have -3 (4) and can multiply freely. -3 * 4 = -12

OpenStudy (anonymous):

to remember what order to do a problem in, just remember PEMDAS. http://www.coolmath.com/prealgebra/05-order-of-operations/images/order-of-operations-05-01.gif

OpenStudy (anonymous):

I know that but this is using the distributive property to write as an equivalent expression

OpenStudy (anonymous):

the distributive property is the property involveing parentheses and multiplication; -12 IS an equivalent expression to -3(7-11), the sign "=" means equal to, or "equivalent".